### v2

Abstract
Strawberry harvesting robots faced persistent challenges such as low integration of visual perception, fruit-gripper misalignment, empty grasping/misgrasp, and strawberry slippage from the gripper due to insufficient gripping force, all of which compromised harvesting stability and efficiency in orchard environments. To overcome these issues, this paper proposed a visual fault diagnosis and self-recovery framework that integrated multi-task perception with corrective control strategies. At the core of this framework was SRR-Net, an end-to-end multi-task perception model that simultaneously performed strawberry detection, segmentation, and ripeness estimation, thereby unifying visual perception with fault diagnosis. Based on this integrated perception, a relative error compensation method based on the simultaneous target-gripper detection was designed to address positional misalignment, correcting deviations when error exceeded the tolerance threshold. To mitigate empty grasping/misgrasp and fruit-slippage faults, an early abort strategy was implemented. A micro-optical camera embedded in the end-effector provided real-time visual feedback, enabling grasp classification during the deflating stage and strawberry slip prediction during snap-off through MobileNet V3-Small classifier and a time-series LSTM classifier. Experiments demonstrated that SRR-Net maintained high perception accuracy. For detection, it achieved a precision of 0.895 and recall of 0.813 on strawberries, and 0.972/0.958 on hands. In segmentation, it yielded a precision of 0.887 and recall of 0.747 for strawberries, and 0.974/0.947 for hands. For ripeness estimation, SRR-Net attained a mean absolute error of 0.035, while simultaneously supporting multi-task perception and sustaining a competitive inference speed of 163.35 FPS. The proposed compensation mechanism reduced the relative errors from 11.52 mm and 5.15 mm to 3.12 mm and 4.11 mm, while maintaining an execution time below 1 s. In addition, the grasp adjustment module responded within 0.5 s, leading to a reduction of about 0.5 s in the deflating stage. During the snap-off stage, when strawberry slipping was predicted, re-inflation followed by a secondary snap-off required 1.88 $\pm$ 0.01 s, achieving a recovery success rate of 81.25%. In cases where slipped was predicted and the operation was terminated, the total execution time was 1.44 $\pm$0.07 s, with slipped status success rate of 88.89%. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, grasp adjustment reduced the cycle time to 4.56 $\pm$ 0.04 s, representing a reduction of approximately 6.7 s, while strawberry slip prediction reduced it to 7.08 $\pm$ 0.10 s, corresponding to a reduction of approximately 4 s. Overall, the proposed framework enhanced both the robustness and efficiency of strawberry harvesting robots, providing a practical and reliable solution for autonomous fruit harvesting.
keywords:
visual fault perception, early diagnosis, relative error compensation, early abort strategy, harvesting robot
1 Introduction
Harvesting robots [hua2025key] [xiao2024review] [2024Detection] have made significant progress in recent years, showing great potential to reduce labor dependence and improve agricultural productivity. However, several mechanical, electrical, and control faults [milecki2023review] [quamar2024review] [rajendran2024towards] still arise during robotic operations, compromising operational stability and continuity. For example, mechanical structural failures [wang2025structural], air leakage in pneumatic end-effectors [rapalo2024design], and fractures in end-joint connectors can impair the normal operation of strawberry harvesting robots. Moreover, the absence of active learning and self-updating mechanisms renders models ineffective in adapting to changing conditions. Failure to respond to abnormal signals also poses operational and safety risks to the robot. Once any single component fails or produces an erroneous output, the entire harvesting process is interrupted. Due to the lack of robust fault diagnosis and self-recovery mechanisms, harvesting robots remain prone to frequent work interruptions, which shortens their effective operation time in the field.
Generally, fruit harvesting with robots involves seven key steps: image acquisition, fruit detection, segmentation and ripeness estimation, instance tracking and localization, motion planning, execution, result evaluation with fault diagnosis, and result recording with optimization. Along this pipeline, potential faults may occur in multiple stages [khalastchi2019fault]. On the perception side, problems such as blurred or occluded cameras, unstable illumination [nejati2020kiwifruit], network/data loss, inaccurate segmentation [shi2025advances] [tantan2025performance], ripeness misclassification [kamat2025multi], ID tracking drift, and depth or calibration errors can lead to detection failures. At the motion control level, incorrect path planning, collision avoidance errors, inverse kinematics failures, or grasp misalignment may disrupt harvesting. Even after successful contact, execution faults such as loose grasps, failed detachments, fruit damage, or gripper jamming remain common. These are further compounded by inaccurate success/failure judgment and inadequate fault recovery, resulting in prolonged downtime. Without robust fault diagnosis and self-recovery mechanisms, current robots are prone to frequent interruptions, severely limiting their effective operating time in orchards.
Figure 1: Images collection devices and environment. (a) HarvestFlex robot with a shading cover; (b) HarvestFlex robot without a shading cover; (c) table-top strawberries
Figure 2: The picking process of strawberry harvesting robot: inflating and approaching, swallowing, deflating, snap off, descending and placing.
In the HarvestFlex system (https://xiong-lab.cn/), as shown in Fig. 1 (a) and (b), a RealSense D455F depth camera (Intel, Realsense, USA) captured real-time RGB-D images, which were processed by YOLOv11 for strawberry detection and segmentation. A checkerboard-based calibration enabled eye-to-hand coordinate transformation, while polynomial regression compensated spatial errors. For sequence planning, a minimal-height sorting algorithm determined harvesting order, and a point-to-point speed pattern-based method ensured efficient move the end-effector. The end-effector was a flexible pneumatic gripper [11231358] that inflated and deflated to gently envelop and detach strawberries. The complete harvesting workflow consisted of six steps: inflating and approaching, swallowing, deflating, snap off, descending and placing, in Fig. 2. However, several issues emerged with our HarvestFlex robot that severely compromised its ability to harvest stably and continuously. These primary challenges were: (1) Low integration of visual perception tasks – the limited fusion of detection, segmentation, and ripeness estimation led to increased model complexity and reduced inference speed, thereby constraining overall system efficiency. (2) Positional inaccuracy between the end-effector and the target fruit – although the eye-to-hand approach offered a stable field of view for dynamic fruit tracking, it remained highly sensitive to hand-eye calibration errors. Even minor deviations could lead to misalignment, empty grasps, unsuccessful detachments, or incorrect placements, thereby disrupting continuous operation. (3) Inadequate gripping and strawberry slippage – although the pneumatic soft gripper reduced fruit damage, its high compliance makes the integration of reliable tactile or slippage sensors challenging. Consequently, slippage often went undetected, resulting in wasted cycle time and, in some cases, unintended fruit release during the snap-off stage. This problem was further compounded by the immaturity of current tactile sensing technologies, which remained expensive, unstable, and impractical for large-scale deployment in orchard environments. Moreover, existing tactile sensor technologies remained immature, with solutions that are typically expensive, unstable, and challenging to deploy reliably in orchard environments.
To mitigate such problems, fault diagnosis and recovery methods were generally classified into two categories: traditional approaches and deep learning-based techniques to deal with electrical, software and control, process-related, and human errors faults [sabry2024review]. Traditionally, various methods were proposed to address robotic faults, such as signal processing techniques [zhang2022vibration], rule-based systems [vukadinovic2024anomaly], model-based approaches [hasan2023model], etc. However, these traditional methods faced limitations in handling complex fault scenarios, particularly in the case of rule-based and threshold-based approaches. With advances in computer hardware, deep learning emerged a new paradigm for fault diagnosis, enabling automatic feature extraction from multi-modal data such as sensor signals, vibrations, acoustics, and images. CNNs were applied to surface defect detection [zhou2025finger], while RNNs and LSTMs were used to analyze time-series data for early fault detection in [hwang2023cooperation]. However, these methods often relied on additional sensors to acquire high-quality multi-modal data, which not only increased system cost and complexity but also complicated deployment in real-world agricultural environments. Furthermore, ensuring temporal and spatial alignment across heterogeneous data sources (e.g., vision, vibration, and tactile signals) remained a nontrivial challenge, as even slight misalignments could degrade fault diagnosis accuracy.
To overcome these limitations, this paper focused on a vision-based fault diagnosis framework for HarvestFlex, thereby avoiding the need for additional multi-modal sensors and the complexities of data alignment. The objective was to address software- and control-related faults that compromised harvesting stability and efficiency. To this end, an end-to-end multi-task perception method SRR-Net [SUN2026111169] was introduced and integrated visual perception with fault object detection while maintaining a lightweight model architecture and low computational load. The method comprised three primary subtasks: object detection, instance segmentation, and ripeness estimation. To align the gripper with the harvesting point, a relative error compensation method based on the simultaneous target-gripper detection was implemented in the coordinate frame of robot arm once the gripper reached the position beneath the target. Then, a corrected harvesting point was generated, guiding the gripper to move beneath the target for a second time to ensure accurate alignment. Furthermore, an early abort strategy was introduced to improve the reliability of the harvesting process. Specifically, a micro-optical camera embedded at the base of the gripper continuously monitors the presence and stability of strawberries. In the deflating stage, the MobileNet V3-Small [9008835] verified whether the fruit had been successfully grasped, while in the snap-off stage, an LSTM classifier estimated the probability of the strawberry slippage from the gripper. If the strawberry was slipping, re-inflation and snap-off actions were triggered. If it had already slipped, an early abort signal was issued to terminate the current cycle. This integrated perception–action framework enabled prompt diagnosis of empty grasps/misgrasp, and slippage, allowing the robotic arm to respond rapidly and maintain stable operation. The main contributions of this paper were as follows.
  • An end-to-end multi-task perception framework, SRR-Net, was introduced to integrate visual perception with fault diagnosis.
  • A relative error compensation method based on the simultaneous target-gripper detection was developed to realign the position of end-effector when approaching the area beneath the picking point.
  • To enhance operational reliability, an early abort strategy was implemented. During the deflating stage, MobileNet V3-Small was used to classify whether the strawberry had been successfully grasped. Subsequently, in the snap-off stage, an LSTM classifier predicted the likelihood of fruit slippage from the gripper, enabling timely corrective actions.
2 Dataset Benchmark
In this paper, FaultData, GraspData, SnapData and SlipData were constructed to support the multi-task vision perception task, grasp adjustment task and enable the time-series LSTM classifier task. All data were collected using the HarvestFlex strawberry-harvesting robot, which includes two robotic arms [chen2024design], two Realsense D455F cameras, as shown in Fig. 1(a–b). The cameras captured RGB-D images at a resolution of 640×480 pixels, with the distance between the lens and the strawberries ranging from 30 to 90 cm. Data were collected in both natural and controlled lighting environments to enhance the robot’s visual adaptability to complex and dynamic orchard conditions.
FaultData For the tasks of detection, segmentation, and ripeness estimation of strawberry and gripper, diverse strawberry formations-such as isolated, overlapping, and occluded fruits-as well as the complete operational sequence of the end-effector during picking images were collected. To ensure robust and continuous operation for multi-task vision perception task, the dataset also incorporated a variety of lighting conditions, weather scenarios, and nighttime environments, laying the foundation for stable, autonomous, all-weather performance. LightStrawberry from [SUN2026111169] also was applied into FaultData. In addition, high-resolution images were captured using a Redmi Note 13 Pro smartphone and an Orbbec Gemini Pro RGB-D camera (Orbbec, Gemini Pro, China) providing enhanced texture and color information. The entire dataset was collected at the Cuihu Factory in Beijing, China, and featured the Fragaria × ananassa ‘Kaorino’ cultivar, as shown in Fig. 1(c). All images were annotated with the polygonal outlines of strawberries and the end-effector using Labelme [torralba2010labelme]. Strawberry with the ripeness of strawberries assigned based on [SUN2026111169] and hand were labeled. The ripeness score was defined within the range [0, 1.1]. The label format for each instance followed the structure: <<cls>> <<ripeness>> <<boundaries>>, where cls represented the object class. The dataset was split into training and validation subsets, consisting of 2954 and 779 images, respectively. Note that the ripeness attribute applied only to strawberries. To ensure a consistent label format, the hand class was assigned alignment flags of 2.
GraspData Using a miniature camera, image data were collected during the deflation stage of the end-effector. The dataset comprised three classes: 389 images of ripe strawberries held in the gripper (label 0), 346 images of empty grasps with no strawberry (label 1), and 388 images of unripe strawberries held in the gripper (label 2). Following a 7:3 stratified split, the training set contained 272, 242, and 272 images for the three classes respectively, while the validation set contained 117, 104, and 116 images.
[…]
SlipData In each frame, key visual features—including the normalized strawberry area, normalized gripper area, normalized background area, and the width, height and center point of the strawberry—were extracted using a fine-tuned SRR-Net based on the SnapData. These features were structured in the format: <<strawberry_area>>, <<gripper_area>>, <<background_area>>, <w><w>, <h><h>, <x><x>, <y><y>, <<label>>. This label indicated whether the strawberry was at risk of slipping from the gripper, with three possible statuses: normal (0), slipping (1), and slipped (2). A sliding window of 5 consecutive frames was used to form each input sequence. For each sequence, a label was assigned based on the next 3 frames immediately following the sequence. This structure enabled the training and validation of a LSTM classifier to estimate the likelihood of strawberry slippage. The sample counts for labels 0, 1, and 2 were 719, 157, and 1962, respectively. Due to the brief duration of the slipping status during the snap-off stage, there is severe class imbalance, particularly for samples labeled as 1 (slipping). To mitigate the issue of sample imbalance, an oversampling method was introduced. Then, each category was split into training and validation sets at a 7:3 ratio. The images from frame 2 to frame 13 were presented, with the frame number and status displayed in the upper left corner of each subfigure in left of Fig. 3. On the right side were the variation curves of the seven feature components throughout the snap-off process, along with their corresponding labels. The horizontal axis represented the frame number, the vertical axis represented the feature values of the corresponding features, and the blue line indicates the labels for the corresponding frames.
Figure 3: Strawberry slip process and status
3 Visual Fault Diagnosis and Self-Recovery Method
3.1 Overview
The HarvestFlex robot has successfully achieved strawberry harvesting in orchard environments. However, several challenges persisted that hindered harvesting performance during continuous operations. The core challenges included: (a) limited integration of visual perception; (b) grasping misalignment or failure resulting from positional offsets; and (c) reliance on human intervention to determine success or failure during the picking process, owing to the absence of automated evaluation and recovery mechanisms, as shown in Fig. 4. A video demo was available at https://1drv.ms/v/c/e68b602748988e28/IQDbhbliJ1nPTqooSIfJg0S0AeDAfkL3JQGtFZfRcWEQImA?e=ISC7Mm.
Figure 4: Harvesting fault and recovery in HarvestFlex
To address the aforementioned issues, a vision-based harvest fault diagnosis and self-recovery architecture was developed. This architecture integrated three key components: an end-to-end multi-task perception network, a relative error compensation method, an early abort strategy based on the empty grasp/misgrasp adjustment and strawberry slip prediction. Specifically, the end-to-end multi-task perception network, SRR-Net, simultaneously performed detection, segmentation, and ripeness estimation. Detection, segmentation, and ripeness estimation were integrated into an end-to-end method. A vision-based positional error compensation method was implemented to correct the $x$- and $y$-axis offsets between the picking point and the gripper center, thereby enhancing localization accuracy. To detect unsuccessful grasps, a micro-optical camera was embedded within the end-effector to monitor strawberry retention during the deflating stage and predict potential slippage risks during the snap-off stage. A MobileNet V3-Small classifier was applied to determine the whether ripe strawberry are present at the end-effector during the deflating stage. An LSTM-based visual time-series prediction model provided real-time forecasts of slippage, enabled a secondary harvesting attempt, the triggering of an early abort signal, or the continuation of subsequent actions.
3.2 End-to-End Multi-Task Perception Method: Detection, Segmentation and Ripeness Estimation
An end-to-end multi-task perception method, SRR-Net, was introduced to simultaneously perform detection, segmentation, and ripeness estimation of both strawberries and gripper. This method integrated low- and high-level feature representations with ripeness and biological attributes to comprehensively perceive strawberries and reduce errors in ripeness estimation. The method employed a lightweight network with shared weights. For strawberries, the perception task distinguished between ripe and unripe fruits by post-processing classification. For the robotic gripper, relative error between the gripper and the picking point was calculated.
This method was built on the general real-time detection and segmentation framework YOLOv11. The architecture consisted of a backbone, neck, and head: the backbone extracted features, the neck fused multi-scale features, and the head performed classification, bounding box regression, instance segmentation, and ripeness estimation. Compared with the original YOLOv11, the primary modification was the addition of a parallel ripeness estimation branch embedded within the head. An overview of the proposed end-to-end multi-task method was shown in Fig. 5.
The ripeness estimation branch consisted of two convolutional layers with a 3×3 kernel, followed by a convolutional layer with a 1×1 kernel and a ReLU activation function. The ReLU constrained the predicted ripeness values within a valid range. To supervise ripeness estimation, the mean absolute error was used as the loss function, scaled by a factor $\lambda$ to balance its influence with the losses of other tasks. The ripeness loss was defined as follows:
$\mathcal{L}_{ripeness}=\frac{\lambda}{N}\sum_{i=1}^{N}|r_{i}-\hat{r}_{i}|$ (1)
where $r_{i}$ was the ground truth ripeness value, $\hat{r}_{i}$ was the predicted ripeness, $N$ was the number of predicted instances, and $\lambda$ was the weighting coefficient.
Figure 5: The architecture of end-to-end multi-task perception method
3.3 Simultaneous Target-Gripper Detection for Relative Error Compensation
Traditional harvesting robots rely on calibration-based methods to transform the object’s location from the camera frame to the robot arm frame. However, these approaches are susceptible to calibration errors and positioning inaccuracies in dynamic environments, exacerbated by system latency, making effective compensation challenging [ge2023three]. In contrast, our proposed method detected the target strawberry and the gripper simultaneously in the same camera view, enabling direct estimation and correction of their relative error within that shared frame. This approach minimized the impact of absolute calibration inaccuracies and system latency, as the offset was computed relative to both elements in the same view, providing a more robust and real-time compensation strategy for improved picking accuracy in the HarvestFlex system.
Figure 6: Simultaneous target-gripper detection for relative error compensation
In the HarvestFlex arm’s workflow, a relative error compensation method was incorporated into the inflating and approaching and swallowing stages of the picking process, in Fig. 6. When the gripper approached the strawberry picking point, the 3D coordinates of both the end-effector and the picking point were obtained in the robot arm coordinate system. Specifically, using SRR-Net, the 2D image coordinates of the strawberry and the end-effector were extracted in the camera coordinate system. By combining these with the corresponding depth information, their 3D positions in the camera coordinate system were reconstructed and then transformed into the robot arm coordinate system through hand–eye calibration. Let $(x_{p},y_{p},z_{p})$ and $(x_{e},y_{e},z_{e})$ denote the coordinates of the picking point and the end-effector in the robot arm frame, respectively. The positional error between the end-effector and the picking point was defined as a 3D offset ($\Delta x$, $\Delta y$, $\Delta z$) , defined as
$\Delta x=x_{p}-x_{e};\Delta y=y_{p}-y_{e};\Delta z=z_{p}-z_{e}$ (2)
Since the snap-off stage required bending the strawberry stem, the vertical position of gripper during the swallowing stage was adjusted to exceed that of the picking point. To simplify computation and reduce processing overhead, only the $x$- and $y$-axis errors were considered, while $z$-axis errors were ignored. Given an error tolerance threshold $T$, when the end-effector reached a position beneath the picking point and the absolute error exceeded $T$, the compensated picking point coordinates $(x_{c},y_{c},z_{c})$ were computed as:
$x_{c}=\begin{cases}x_{p}+k_{x}\Delta x&\text{if }|\Delta x|>T\\ x_{p}&\text{otherwise}\end{cases}$ (3)
$y_{c}=\begin{cases}y_{p}+k_{y}\Delta y&\text{if }|\Delta y|>T\\ y_{p}&\text{otherwise}\end{cases}$ (4)
$z_{c}=z_{p}$ (5)
Here, $T$ denoted the maximum tolerance of positional error for each axis, set to 10 mm. $k_{x}$ and $k_{y}$ referred to the scale factors of error. The robotic arm was then moved to the compensated point for strawberry grasping.
3.4 Early Aborting Robotic Picking upon Grasp Adjustment and Slip Prediction
Empty grasp and misgrasp represented common failure modes in robotic strawberry harvesting, particularly during the deflating and snap-off stages. During the deflating stage, empty grasps occurred when the strawberry was not successfully enclosed by the end-effector, primarily due to localization errors or unexpected fruit motion. In the snap-off stage, strawberry slippage occurred when the picked strawberry unintentionally slipped, often due to a non-optimal end-effector pose or inadequate gripping force. To diagnose such faults effectively, a simple yet efficient method involved the installation of a miniature camera (JTS302, JiuTan, China) with a 3 mm diameter at the bottom of the end-effector, as shown in Fig. 7. The micro-optical camera directly captured real-time images of the gripper area, enabling visual monitoring of strawberry presence.
Figure 7: Gripper with a micro-optical camera. (left) Top-view and (right) Side-view
Once the deflating stage began, a MobileNet V3-Small was used to determine whether a strawberry was present in the gripper, as shown in Fig. 8. If the classification result is 0 (strawberry correctly grasped), the robotic arm continued executing the subsequent picking actions. If the result is 1 (empty grasp) or 2 (misgrasp of unripe strawberry), the manipulator re-inflated the gripper, executed a descending motion, skipped the snapping and placing stages, and directly performed the reset motion to terminate the current picking cycle. Instead, it immediately selected a new target strawberry and proceeded to the next picking cycle. The process flow, indicated by the blue line, was illustrated in Fig. 4.
Figure 8: Empty grasp and early abort
For the slippage scenario, a time-series LSTM classifier was developed to predict whether the strawberry would slip from the end-effector during the snap-off stage. The LSTM classifier architecture was summarized in Table 1. The input layer of LSTM classifier had 7 features based a finetuned SRR-Net wtih strawberry and background classes. The LSTM layer consisted of 5 layers, a hidden size of 64, and a dropout of 0.2. This was followed by a Dropout layer with a rate of 0.3, then a Linear layer mapping from 64 to 3 units, and a Softmax layer applied along dimension 1. The output layer produced 3 classes.
Table 1: LSTM classifier architecture
Layer Number
input 7
LSTM layers=5
hidden=64
dropout=0.2
Dropout 0.3
Linear (64, 3)
Softmax dim=1
output 3
The input of the LSTM classifier consisted of a sequence of features from the past 5 frames, represented as <<normalized strawberry area, normalized gripper area, normalized background area, $w$, $h$, $x$, $y>$. The normalized strawberry area, gripper area, and background area referred to the proportion of pixels in the image corresponding to the strawberry, the gripper, and the background, respectively. $w$ and $h$ represented the width and height of the strawberry. The model outputted the probability of slippage over the following three frames. If the probability fell below the minimum threshold, the strawberry remained attached to the end-effector after the snap-off motion, indicating a successful pick. Conversely, if the probability was equal to or exceeded the maximum threshold, the strawberry had slipped, indicating a failure in the snap-off stage. When the probability fell between the minimum and maximum thresholds, the strawberry was in the process of slipping from the gripper.
Figure 9: Strawberry slip prediction and self-recovery
To enhance reliability, a time-stability rule was implemented. If slipped was predicted in two consecutive instances, an early abort signal was triggered, the placing stage was skipped, and the next picking cycle began. If two consecutive frames were slipping, the end-effector retracted, reinflated, and re-performed the snap-off action. Otherwise, the robotic arm performed the subsequent actions.
3.5 Paradigm of HarvestFlex Visual Fault Diagnosis and Self-Recovery
In HarvestFlex, visual fault diagnosis and self-recovery proceeded through the following steps, in Fig. 10. First, RGB and depth images were captured by a RealSense camera. The RGB images were processed by an end-to-end multi-task perception network to detect and segment strawberries and the gripper, and to estimate strawberry ripeness. Using a predefined ripeness threshold, fruits were classified as ripe or unripe, and only ripe strawberries were tracked in real time. Next, the camera coordinates of both tracked ripe and untracked unripe strawberries were transformed into the robot arm coordinate via hand–eye calibration. The end-effector’s coordinates, based on the camera coordinate system, were also extracted, tracked, and transformed to prevent track ID jumps. The system then planned the picking sequence for ripe strawberries and computed the robot arm motion trajectory to reach the (compensated) picking points. The gripper harvested each strawberry in sequence until no ripe fruits remained, at which point the mobile platform repositioned.
Figure 10: Paradigm of HarvestFlex visual fault diagnosis and self-recovery
The relative positional error compensation method, incorporating visual feedback SRR-Net, and the early abort strategy based on MobileNet V3-Small and the LSTM-based strawberry slip prediction were implemented as follows. The gripper then inflated and approached a location directly beneath the target strawberry picking point. At this position, the relative error between the gripper and the actual strawberry position was calculated. If the error exceeded a predefined tolerance threshold, a compensated picking point was generated to correct the deviation before proceeding.
Once aligned, the gripper executed swallowing and deflating motions to enclose the strawberry. During this process, a micro-optical camera mounted on the gripper captured images, which were used to determine whether the strawberry was successfully gripped based on MobileNet V3-Small classifier. When no strawberry or an unripe strawberry was detected in the gripper for two consecutive frames, the robot triggered re-inflation, descent, and homing actions to begin a new picking cycle. Conversely, it continuously performed the snap-off action while running a time-series LSTM classifier to predict strawberry normal.
To improve prediction stability, a time-stability rule was applied. When two consecutive slipped signals were triggered, the robot aborted the snap-off stage, skipped the placing stage, and returned to its homing position, as illustrated by the red lines in Fig. 10. When two consecutive normal signals were triggered, the robot continued the snap-off motion, as illustrated by the black line. Meanwhile, if two consecutive slipping signals were triggered, the gripper was inflated and deflated to re-grasp the strawberry during the deflating and snap-off actions, followed by the descending, placing, and homing steps, as illustrated by the orange line.
4 Experiments
4.1 Evaluation Metrics
For the detection and segmentation tasks, precision and recall were used to evaluate performance, following the definitions in [10028728]. For strawberry ripeness estimation, the mean absolute error (MAE) was adopted as the evaluation metric, defined as:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$ (6)
where $y_{i}$ denoted the ground-truth ripeness value, $\hat{y}_{i}$ was the predicted ripeness, and $N$ was the total number of samples. Additionally, the inference speed was assessed by measuring frames per second (FPS) on the FaultData validation set with resolution of 640×480.
For the relative error between the end-effector and the picked strawberry instance, the visual relative errors along the $x$-, $y$-, and $z$- axes under the robotic arm coordinate system were denoted as $\Delta x$, $\Delta y$, and $\Delta z$, respectively. Due to the bending motion required during the snap-off stage, the end-effector’s ascent height could exceed that of the target point. Therefore, the error calculation was simplified from the $x$-, $y$-, and $z$-axes to only the $x$-, and $y$- axes. Similarly, the corresponding physical errors along the $x$-, and $y$-axes were denoted as $\Delta x_{w}$ and $\Delta y_{w}$. The physical error after compensation, denoted as $E_{x}$ and $E_{y}$, was measured to validate the effectiveness of the relative error compensation method. Finally, mean value $\bar{e}$ was used to evaluate the accuracy of the relative error and physical error before compensation and for each coordinate axis and can be formulated as:
$\bar{e}=\frac{1}{n}\sum_{1}^{n}|e|,\text{ where }e=\{\Delta x,\Delta y,\Delta x_{w},\Delta y_{w},E_{x},E_{y}\}$ (7)
In addition, FPS (frames per second) was used to evaluate the inference speed of the model.
4.2 Implementation Details
To evaluate the effectiveness of the proposed method, experiments were conducted on the FaultData, GraspData, SnapData, and SlipData benchmarks. The end-to-end multi-task perception framework was first trained and validated on the training and validation subsets of FaultData. During this evaluation, the relative error of the strawberry and the gripper beneath the picking point were measured, along with the physical error before compensation, the relative error before compensation, and the physical error after compensation, to assess picking accuracy. Furthermore, the end-to-end multi-task perception method was fine-tuned using the images from the mini-camera. The outputs of the multi-task perception model were then applied to generate the SlipData for LSTM classifier. The model was trained using an NVIDIA RTX 4090 GPU.
4.3 Results of Detection, Segmentation, Ripeness Estimation
SRR-Net, which integrated multi-task learning for both strawberry and hand classes, was evaluated on the FaultData dataset for detection, segmentation, and ripeness estimation in Table. 2. For bounding box prediction, SRR-Net achieved a precision of 0.895, recall of 0.813, mAP@50 of 0.884, and mAP@50-95 of 0.633 for strawberries, while for the hand class, the corresponding values were 0.972, 0.958, 0.977, and 0.788, respectively. In segmentation, SRR-Net reached a precision of 0.887, recall of 0.747, mAP@50 of 0.829, and mAP@50-95 of 0.448 for strawberries, and 0.974, 0.947, 0.964, and 0.655 for hands.
Compared with YOLOv11 and YOLOv11-seg, SRR-Net demonstrated comparable detection and segmentation performance across both classes, with minor variations in individual metrics. In detection, its recall for strawberries was slightly higher than YOLOv11, while the mAP values remained similar to YOLOv11-seg. In segmentation, SRR-Net showed a modest gain in recall over YOLOv11-seg, while maintaining comparable precision. These results indicate a balanced trade-off between precision and recall, with the segmentation branch preserving high precision and competitive recall, enabling reliable identification of strawberries and hands.
In ripeness estimation, SRR-Net achieved a mean absolute error of 0.035, highlighting its advantage in accurately assessing strawberry ripeness levels. Regarding inference speed, SRR-Net reached 163.35 FPS, remaining close to the speed of YOLOv11-seg (167.89 FPS), demonstrating that the added ripeness estimation branch does not compromise real-time performance.
Overall, these results suggested that SRR-Net maintained detection and segmentation performance on par with established baselines, while also demonstrating solid capability in ripeness estimation accuracy and inference speed. As an end-to-end multi-task visual perception framework, it effectively integrated detection, segmentation, and ripeness estimation, achieving a well-balanced performance across all tasks.
Table 2: The detection, segmentation, and ripeness estimation results on FaultData and and SnapData.
Method Class Box Mask MAE FPS
P R mAP@50 mAP@50-95 P R mAP@50 mAP@50-95
FaultData
YOLOv11 strawberry 0.915 0.792 0.884 0.633 - - - - - 201.10
hand 0.966 0.957 0.977 0.793 - - - - -
YOLOv11-seg strawberry 0.892 0.812 0.884 0.663 0.864 0.753 0.824 0.442 - 167.89
hand 0.968 0.957 0.976 0.791 0.971 0.95 0.975 0.653 -
SRR-Net strawberry 0.895 0.813 0.884 0.633 0.887 0.747 0.829 0.448 0.035 163.35
hand 0.972 0.958 0.977 0.788 0.974 0.947 0.964 0.655 -
SnapData
SRR-Net strawberry 0.991 0.99 0.992 0.98 0.991 0.99 0.994 0.984 - -
background 0.928 0.911 0.968 0.863 0.928 0.911 0.963 0.859 -
To more intuitively observe the performance of SRR-Net, the visualization results of SRR-Net, YOLOv11, and YOLOv11-seg on the FaultData dataset are presented in Fig. 11. In Fig. 11 (a) and (b), strawberries and end-effectors were represented with bounding boxes, with the target class label and confidence score displayed above each box. The confidence threshold was set to 0.5. The ripeness value of SRR-Net was displayed between the target class label and the confidence score. For the same strawberry instance, SRR-Net produced higher confidence scores than YOLOv11 and YOLOv11-seg. In terms of ripeness estimation, ripeness estimation of SRR-Net closely matched the true ripeness of each strawberry. Overall, the visual comparisons illustrated that SRR-Net delivered superior performance compared with YOLOv11 and YOLOv11-seg.
(a) YOLOv11
(b) YOLOv11-seg
(c) SRR-Net
Figure 11: Visualization results
4.4 Evaluation of Relative Error Compensation
To validate the effectiveness of relative error compensation, the 3D coordinates of the strawberry picking point ($x_{s}$, $y_{s}$, $z_{s}$), the gripper positioned beneath the picking point ($x_{e}$, $y_{e}$, $z_{e}$), the relative error before compensation ($\Delta x$, $\Delta y$), the ground-truth error before compensation ($\Delta x_{w}$, $\Delta y_{w}$) , the compensated picking point ($x_{ce}$, $y_{ce}$, $z_{ce}$), and the relative error after compensation ($E_{x}$, $E_{y}$) in the robot arm coordinate system were measured in Table 3. Due to the requirement of snap-off, the rise distance in the swallowing stage exceeded the height of the strawberry. To simplify the experiments and reduce computational overhead, the relative error along the $z$-axis was not considered.
From Table 3, the visual and ground-truth errors before compensation were compared, and the ground-truth errors after compensation were recorded. For example, in the first row, the visual relative errors on the $x$- and $y$- axis were 22 mm and -4 mm, while the corresponding ground-truth errors were 17.3 mm and 4.2 mm. The absolute differences between visual and ground-truth errors were therefore 4.7 mm and 8.2 mm. After compensation, the physical errors of the $x$- and $y$- axis were reduced to 1.5 mm and 6.3 mm, both below the defined threshold.
Before compensation, the mean relative errors were 14.07 mm on the $x$-axis and 8.64 mm on the $y$-axis, while the mean physical errors were 11.52 mm and 5.15 mm, respectively—indicating that visual estimation tended to overestimate the actual physical errors. With the proposed compensation method, the mean physical error between the strawberry point and the gripper were further reduced to 3.12 mm ($x$-axis) and 4.11 mm ($y$-axis). The slightly larger residual $y$-axis error was likely due to the difficulty in achieving high-precision motor control over very small movement ranges. Despite this phenomenon, the relative error compensation method demonstrated a clear advantage in aligning the gripper with the target and in reducing grasping errors.
Table 3: Relative errors between the end-effector and the picking point in the arm coordinate system (unit: $m$). $(x_{s},y_{s},z_{s})$ denoted the coordinates of the initial picking point, and $(x_{e},y_{e},z_{e})$ represented the end-effector position when it moved to a location beneath the strawberry. The visual relative error between the vision-based picking point and the end-effector was defined as $\Delta x$ and $\Delta y$, while $\Delta x_{w}$ and $\Delta y_{w}$ denoted the corresponding positional error in the physical world. After applying the compensation strategy, the updated picking point was represented as $(x_{ce},y_{ce},z_{ce})$. The relative error between the compensated picking point and the end-effector in the physical world was denoted as $E_{x}$ and $E_{y}$. The final row showed the computed mean absolute error.
$x_{s}$ $y_{s}$ $z_{s}$ $x_{e}$ $y_{e}$ $z_{e}$ $\Delta x$ $\Delta y$ $\Delta x_{w}$ $\Delta y_{w}$ $x_{ce}$ $y_{ce}$ $z_{ce}$ $E_{x}$ $E_{y}$
709 221 706 686 225 647 22 -4 17.3 4.2 732 219 706 1.5 6.3
464 232 710 450 249 652 14 -17 8.3 -12.7 479 223 710 5.0 -8.2
377 245 693 362 255 626 15 -10 7.2 -2.5 392 239 693 2.6 3.6
700 222 699 681 224 638 18 -2 15.7 3.4 719 221 699 0 2.8
532 234 710 518 245 649 14 -11 10.6 -3.6 546 228 710 2.7 -2.9
711 244 711 692 249 651 18 -5 9.2 -5.1 730 241 711 0 2.2
320 244 692 316 256 628 4 -11 3.6 -2.6 324 239 692 0 2.1
468 236 703 460 245 653 8 -8 12.8 -1.3 - - - - -
652 235 712 631 239 651 20 -3 20.4 6.2 673 233 712 -4.4 -2
816 220 699 793 223 642 23 -2 17.4 6.8 839 219 699 -6.4 5.1
393 238 706 386 245 643 7 -7 6.9 3.0 - - - - -
734 222 699 712 224 638 22 -1 11.8 5.0 757 221 699 -5.3 3.9
445 230 711 436 249 646 8 -19 9.9 -9.5 454 220 711 -4.3 6.2
629 234 721 611 244 659 17 -10 17.5 -2.6 647 228 721 5.9 6.3
299 246 693 295 259 631 4 -12 10 -7.9 304 240 693 -2.4 1.2
453 233 701 443 251 637 10.6 -18.5 10.4 -7.5 464 224 701 7.1 -1.2
752 213 708 729 218 645 22 -4 13.4 6.9 774 221 708 0 4.3
667 215 708 645 222 645 22 -7 14.2 0 690 211 708 -3.8 -6.0
307 246 693 305 252 633 1.71 -6.3 5.7 0 - - - - -
467 232 700 456 248 634 11 -15 8.1 -12.1 479 224 700 -1.6 5.6
- - - - - - 14.07 8.64 11.52 5.15 - - - 3.12 4.11
4.5 Evaluation of Empty Grasp or Misgrasp
MobileNet-V3 Small was trained and validated based on the GraspData. The classification results were presented in Table 4. Class 0, 1, and 2 represent the presence of strawberries, the without strawberry, and the unripe strawberry in the end-effector, respectively. From the Table 4, it can be observed that the precision, recall, and F1-score for all classes are all 1.00, indicating that the model achieves optimal classification performance for each category and closely matches the ground-truth labels. The classification results for all categories were 1, mainly because the distinguishing features between missed grasps and incorrect grasps are relatively prominent. Overall, the classification results were excellent. To observe the classification results of MobileNet-V3 Small, the classification outputs of some images were visualized in Fig. 12. The upper-left corner of each image was annotated with the predicted class and the corresponding probability. From Fig. 12, it can be concluded that MobileNet-V3 Small achieved good classification performance.
Table 4: Results of empty grasp or misgrasp
Classes Precision Recall F1-score
0 1.00 1.00 1.00
1 1.00 1.00 1.00
2 1.00 1.00 1.00
Based on the classification results of MobileNet-V3 Small, different responses were triggered on the robotic arm. When the classification result was 0, the robotic arm continued to execute the subsequent operations. When the classification result was 1 or 2, the end-effector performed secondary inflation to prevent damage to the strawberry, then skipped the snap-off and placing stages and directly moved the end-effector downward to a position below the strawberry, thereby ending the current picking cycle.
(a) Strawberry in the end-effector
(b) Strawberry without in the end-effector
(c) Misgrasp of unripe strawberry
Figure 12: Visualization of empty grasp or misgrasp
4.6 Evaluation of Slip prediciton
Before performing LSTM classification in the snap-off stage, SRR-Net was first fine-tuned based on SnapData. These images contained two classes—strawberry and background—which were used to count the number of pixels belonging to the strawberry and background regions during the snap-off stage. The results of SRR-Net on SnapData were presented in Table 2. The precision and recall of the segmentation task reached 0.991 and 0.990 for strawberries, and 0.928 and 0.911 for the background, respectively, indicating that the model achieves highly accurate segmentation performance.
Table 5: LSTM classifier result on the SlipData
Class Precision Recall F1-socre
0 0.94 0.89 0.92
1 0.86 0.94 0.90
2 0.99 0.94 0.96
The normalized strawberry area, normalized background area, normalized hand area, width and height of strawberry, center point of strawberry as input of LSTM classifier to train and validate using SlipData. The results of the LSTM classifier on the validation set of SlipData were presented in Table. 5, where class 0, 1 and 2 represented strawberries remaining in the gripper, slipping and slipped, respectively. As shown in Table. 5, the LSTM classifier achieved high accuracy despite the imbalance among class samples.
Visualization results of the LSTM classifier on robots in the orchard were presented in Fig. 13. The left panel displayed frames 3-14 with the prediction probability, current strawberry status, the prediction probability of three status, and final strawberry status based on time-stability, while the right panel illustrated the LSTM prediction probabilities. The horizontal axis represented the frame index starting from 0, and the vertical axis indicated the predicted probability of strawberry slippage within the next three frames. If the total number of frames did not reach 5, fewer than 10 frames were displayed as labels in the top-left corner of the images. The LSTM classifier predicted a slipping status at frames 7, 11, and 12, with corresponding probabilities of 0.62, 0.70, 0.82. Based on the time-stability criterion, secondary inflating and snap-off were triggered at frame 12. The probability curve of strawberry slip prediction was plotted for intuitive observation, as shown in the right panel of Fig. 13. Similarly, a probability curve of strawberry slippage was plotted for intuitive observation in Fig. 13 right.
Figure 13: LSTM classifier results. (Left) Video frames 3-14; (Right) Predicted probabilities from the LSTM model. Note that frames were counted starting from 0.
4.7 Operational Efficiency Analysis of the Early Abort Strategy
[…]
Table 6: Harvesting time of robot arm (unit: s). – indicated that the execution time at the corresponding stage under a certain fault was identical to the original (baseline) time. $\times$ indicated that the corresponding stage was not executed.
Action Original Empty Grasp Misgrasp Slipped Compensation
Inflating & approaching 1.25 $\pm$ 0.01
Compensation $\times$ $\times$ $\times$ $\times$ 0.71 $\pm$ 0.07
Swallowing 0.73 $\pm$ 0.00
[…]
Descending 0.98 $\pm$ 0.08
Placing 4.36 $\pm$ 0.01 $\times$ $\times$ $\times$
Homing 1.88 $\pm$ 0.00
[…]
For grasp adjustment, empty grasp times and misgrasp time were shown in Fig. 15. The response times for empty grasp and misgrasp were 0.42 $\pm$ 0.04 and 0.39 $\pm$ 0.03, respectively. Compared with original deflating stage, grasp adjustment can save approximately 0.5 s. The snap-off and placing stages were skipped, which saved time in these two phases. Compared with baseline method, the mitigation of grasp adjustment reduced the harvesting cycle time by approximately 4 seconds, thereby improving overall harvesting efficiency in Table. 6.
Figure 14: Relative Error Compensation Time
Figure 15: Response time of grasp adjustment during the deflating stage (unit: s).
For strawberry slip prediction, the harvesting time for strawberries with normal status was 1.03 $\pm$ 0.00 s. For strawberry slipping that triggered re-inflating and re-snap off, the execution time was 1.81 $\pm$ 0.07 s. For slipped, the total time was 1.44 $\pm$ 0.07 s. It should be noted that the time recorded for the confirmed slippage state includes the duration of re-inflation and re-bending attempts, followed by the final detachment of the strawberry. During the snap-off stage, if the strawberry was detected to be slipping, the completion time for secondary inflation, grasping signal generation, and manipulator response was 1.81 $\pm$ 0.07 s. When the strawberry had already slipped off, the manipulator response time decreased to 1.44 $\pm$ 0.07 s. With early termination enabled, the total execution time of the harvesting cycle was reduced to 7.08 $\pm$ 0.10 s. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, the proposed strategy reduced the harvesting cycle time by approximately 4 s. The picking cycle time increased to 12.51 $\pm$ 0.09 s after integrating relative error compensation for grasping and slip prediction, compared with the original cycle time of 11.21 $\pm$ 0.15 s, representing an increase of approximately 1.3 s.
Table 7: Robot arm response time of strawberry slip prediction (unit: s).
Strawberry Status Normal Slipping Slipped
Time 1.03 $\pm$ 0.00 1.81 $\pm$ 0.07 1.44 $\pm$0.07
Furthermore, the success rates of grasp verification and strawberry slippage prediction were assessed through repeated trials based on the robotic arm’s response. As shown in Table. 8, the system achieves an approximately 90.00% success rate in identifying empty grasp or grasp of unripe strawberry, thereby ensuring appropriate corrective actions on the robot arm. For strawberry slip prediction during the snap-off stage, the success rate for strawberries in a normal state reached 96.88%. For slipping, the success rate of secondary grasping was 81.25%, while slipped was 88.89%. The relatively lower success rate for initiating secondary inflation and snap-off during snap-off was mainly attributed to the short duration of the slip event, which requires rapid prediction within a limited time window. In addition, the performance was influenced by the initial grasping posture of the end-effector during the first contact with the strawberry.
Table 8: Statistics on robotic arm response success rate for repeated grasp judgments and slip predictions
Result Grasp Adjustment Strawberry slip prediction
Empty Grasp Misgrasp Normal Slipping Slipped
Success 27 30 31 26 32
Failure 3 3 1 6 4
Success Rate 90.00% 90.91% 96.88% 81.25% 88.89%
5 Discussion
Vision-based fault diagnosis and self-recovery offered an effective means of enhancing the stability and efficiency of strawberry-harvesting robots. During the inflating and approaching stage, the positional relationship between the gripper and the harvesting point served as a visual indicator of accumulated errors caused by fruit recognition, localization, hand–eye calibration, and inverse kinematics. Without requiring an additional camera, relative errors based on the simultaneous target-gripper detection were computed to align the strawberry with the end-effector. For early abort feedback during the deflating and snap-off stages, a micro-optical camera was embedded in the end-effector to detect and predict the probability of strawberry slippage in the gripper. This approach compensated for the absence of force feedback in the flexible pneumatic gripper, whose deformation during inflation and deflation rendered conventional force sensing impractical. Mechanical failures, inverse kinematics errors, and other control-related faults were not addressed in this work.
Additionally, self-learning, continual learning and adaptive evolutionary perception, planning and decision-making can be applied to reduce dataset dependency and improve the generalization and robustness of the strawberry-harvesting robot. With the rapid development of embodied artificial intelligence and robotic agents, a new wave of end-to-end, multi-modal, large-model-based perception, self-planning and self-decision frameworks is emerging in the robotics domain. These advances offer valuable insights for building strawberry harvesting robots and enabling collaborative operations among multiple robots. In the future, developing end-to-end active learning and continuously evolving methods is expected to become a major research trend.
6 Conclusion
This paper proposed a visual-based fault diagnosis and self-recovery system to address gripper offset and strawberry slippage during harvesting. An end-to-end multi-task perception network with a shared weight backbone and neck was developed to reduce computational overhead. Without the need for an additional external camera, relative error compensation method based on the simultaneous target-gripper detection was estimated, enabling compensation actions to mitigate cumulative errors. To monitor the strawberry’s status in the gripper, a miniature optical camera was embedded at its base. A MobileNet V3-Small classifier was adapted to detect strawberries within the gripper and trigger early abort signal when necessary during the deflating stage based on the GraspData. For strawberry slippage during the snap-off stage, a novel dataset, SnapData, comprising strawberry and background classes was introduced, and the perception network was extended and fine-tuned to segment strawberries, the gripper, and background. Width, height and center point features were combined with a time-series LSTM classifier to predict strawberry slippage from the gripper on the SlipData, enabling early-abort picking actions and improving harvesting efficiency. Experimental results demonstrated that the perception network achieved high accuracy, with a mean absolute error of 0.035 for ripeness estimation. The compensation mechanism reduced relative errors to 3.12 mm and 4.11 mm, with a required execution time of less than 1 s. The response time of the grasp adjustment was approximately 0.4 s, resulting in a time saving of about 0.5 s during the deflating stage. For strawberry slipping status during the snap-off stage, re-inflation followed by a second snap-off required 1.81 $\pm$ 0.07 s and achieved a recovery success rate of 81.25%. When the operation was terminated due to slipped, the execution time was 1.44 $\pm$0.07 s, achieving a detection success rate of 88.89%. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, grasp adjustment reduced the cycle time by approximately 4 s, and strawberry slip prediction reduced it by approximately 4 s as well. Overall, the proposed system effectively enhanced the robustness and efficiency of strawberry harvesting, providing a practical solution for reliable autonomous fruit picking. A demonstration video of visual diagnosis and self-recovery is available at https://1drv.ms/v/c/e68b602748988e28/IQAxE3rnaIYoTY9jHLkIULlnAZYotL9PTijSNry9i0aMnw4?e=HJhCL4.
Declarations
[…]
Funding This work was supported by the Haidian District Bureau of Agriculture and Rural Affairs, the Beijing Academy of Agriculture and Forestry Sciences (BAAFS) Innovation Ability Project (KJCX20240321), the Outstanding Youth Foundation of BAAFS (YKPY2025007), the BAAFS Talent Recruitment Program and the National Natural Science Foundation of China (NSFC) Excellent Young Scientists Fund (overseas).
[…]
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
Data availability The data will be made publicly available.
[…]
Code availability The code will be made publicly available.
[…]
Author contribution Meili Sun: Methodology, Software, Validation, Writing - original draft. Chunjiang Zhao: Investigation, Funding acquisition, Writing - review editing. Lichao Yang: Data curation, Methodology, Software, Visualization. Hao Liu: Data curation, Methodology, Software. Shimin Hu: Data curation, Software, Visualization. Ya Xiong: Conceptualization, Methodology, Investigation, Funding acquisition, Writing - review editing.